OpenStudy (anonymous):

How do I balance the equation: CH4+O2= CO2+ H2O?

OpenStudy (anonymous):

This is how I learned it: In order for the equation to be balanced, the number of elements have to be the same on both sides of the equal sign. Original Equation: CH4 + O2 = CO2 + H2O In the original equation, there is one Carbon, four hydrogen, and two oxygen on the left side of the equal sign. Then there is one carbon, three oxygen, and two hydrogen on the right side. 1 = 1 Carbon is balanced and you don't need to do anything to it. 4 is not equal to 2, therefore the hydrogen needs to be balanced. 2 is not equal to 3, therefore the oxygen needs to be balanced. Add a 2 in front of O2 and in front of H2O. Balanced equation: CH4 + 2 O2 = CO2 + 2 H2O In the balanced equation, there is one carbon, four hydrogen, and four oxygen on the left side. On the right side, there is one carbon, four oxygen, and four hydrogen. 1 = 1 4 = 4 4 = 4 Hope this helps.

OpenStudy (anonymous):

Yes, thank you.